OpenStudy (anonymous):

If f(x) = x + 2, find f(3).

OpenStudy (anonymous):

replace all x's by 3 and then evaluate the right hand side

OpenStudy (anonymous):

?

OpenStudy (anonymous):

replace the x with 3's, as such:\[f(x)=x+2~~~\rightarrow~~~f(3)=3+2\]now just solve:\[f(3)=3+2\]

OpenStudy (anonymous):

do you remember substitution? f(3) is the same as saying x=3 for the equation f(x)

OpenStudy (anonymous):

I dont know how to do it still...

OpenStudy (anonymous):

what part is confusing to you?

OpenStudy (anonymous):

How to get the answer

OpenStudy (anonymous):

i dont know how to solve it even tho u did most of it like how would i finish it

OpenStudy (anonymous):

f(3) means that you are finding the answer of the function if the value of x was 3....so you substitute 3 for the x and write it as f(3) = 3 + 2 and solve it....you would get f(3) = 5 and that means that when x = 3, the value of the function is 5

OpenStudy (anonymous):

does that help @darkbbn ? :)

OpenStudy (anonymous):

oh ok i get it u added the 3 and the 2 :D

OpenStudy (anonymous):

yes because we substituted 3 for the x :)